OpenStudy (anonymous):

Angela made the trip by bus in 16 hours and the return trip by car in 20 hours. A bus traveled 10 miles per hour faster than the car. How far was the trip? (Use the formula D=rt [Distance = rate x time])

OpenStudy (anonymous):

Do you have any idea what to do?

OpenStudy (anonymous):

No I don't lol

OpenStudy (anonymous):

That's ok! :) We have to set two equations: one for the car and one for the bus.

OpenStudy (anonymous):

FOR THE CAR: time: 20 hours rate: x FOR THE BUS: time: 16 hours rate: 10+x \(\text{(since it is 10 mph faster than the car)}\)

OpenStudy (anonymous):

we know that d=rt and e have to find the distance.

OpenStudy (anonymous):

EQUATION FOR THE CAR: d=20x EQUATION FOR THE BUS: d=16 (10+x)

OpenStudy (anonymous):

Are you getting it?

OpenStudy (anonymous):

So far yes

OpenStudy (anonymous):

Okay cool! So now, equate the two equation since both of them have d=??? So... \(\huge 16(10 + x)=20x\) Now solve for x.

OpenStudy (anonymous):

Is it x = 8?

OpenStudy (anonymous):

\(\huge 160+16x=20x\)

OpenStudy (anonymous):

\(\huge 160=4x\) What is \(160 \div 4\)?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

Good! So we now know that x=40. Now, substitute it to either of the two equations: EQUATION FOR THE CAR: d=20x EQUATION FOR THE BUS: d=16 (10+x)

OpenStudy (anonymous):

800

OpenStudy (anonymous):

So the distance is 800. :)

OpenStudy (anonymous):

Thank you! Can you help me with another problem lol

OpenStudy (anonymous):

I can try lol

OpenStudy (anonymous):

It's not a distance problem, it's a percent problem

OpenStudy (anonymous):

Forty percent of the signs are circular. If 3600 signs are noncircular, how many signs are there in all?